# Integrating Open-World Shared Control in Immersive Avatars

### v1

@@ VI Experiments
TABLE I: Differences between each interface across all tasks. $*{}^{*}$, ${}^{**}$, and ${}^{***}$ denote $p\leq 0.05$, $p\leq 0.01$, and $p\leq 0.001$ respectively.
Condition Success (%) ()($\uparrow$)( ↑ ) Time (s) ()($\downarrow$)( ↓ ) Usability ()($\uparrow$)( ↑ ) Workload ()($\downarrow$)( ↓ ) Presence ()($\uparrow$)( ↑ )
@@
Friedman p-value ${}^{***}$0.0005 ${}^{**}$0.0060 ${}^{**}$0.0065 ${}^{***}$0.0007 0.6004
Post-hoc p-value DT vs. MM ${}^{**}$0.0066 0.0611 ${}^{**}$0.0015 ${}^{**}$0.0053 0.1308
DT vs. PM ${}^{***}$0.0004 $*{}^{*}$0.0115 ${}^{**}$0.0061 ${}^{***}$0.0004 0.5202
@@
Human subjects studies were conducted to evaluate differences between the DT, MM, and PM interfaces. We formulated the following a priori hypotheses about the system:
@@
We recruited 20 student participants from the University of Illinois at Urbana-Champaign campus, 19 of whom completed the entire procedure. One participant requested to end the experiment during training due to nausea. Of the 19 participants, 11 were male, 7 were female, and one preferred not to say. Subjects were of age 19–32 (mean: 24) and self-reported their familiarity with robotics and controlling robots on average as 5.4 and 4.4 on a 7-point Likert scale [27] respectively. None of the subjects had used TRINA before.
@@ VI-1 Basic Training
Subjects were trained to use the direct teleoperation interface and were introduced to several possible fault states. For example, if excessive force was applied to the arm, the subject would momentarily lose control of it. Subjects were given suggestions about how to resolve each of these faults. The assistive functionalities were then demonstrated using the manual and predictive menus.
@@ VI-2 Task Introduction
Subjects were shown the three testing tasks and completed the tasks in-person to familiarize themselves with the specific features of the target objects. A researcher explained how task completion would be graded, and that subjects should try to complete tasks as quickly as possible with 5 min at most for each task. For the jar, the task was completed when the lid no longer was touching the jar body. For the whiteboard, the required writing was split into 19 segments and credit was given for each completed segment. For the plug, the task was completed when the subject had fully inserted the plug into the target socket so that the face of the plug was touching the socket.
@@ VI-3 Training Tasks
Subjects were coached through using the manual and predictive menus on two training tasks. In the first task, a researcher handed TRINA a capped Expo marker, and the subject had to use TRINA to insert the tip of the marker into a square hole, shown in 5(a). Subjects were told to use the menu to snap to the plane of the hole and turn off all rotational DoFs before inserting the marker into the hole. In the second task, subjects had to grasp and turn a dial shown in 5(b) for three full rotations. They were instructed to first snap to the circle of the dial, disable all but the x and roll DoFs to grasp the dial, and finally have only roll enabled to turn the dial.
(a) Peg-in-hole training rig.
(b) Dial training rig.
Figure 6: Training rigs for the operator. [Best viewed in color.]
@@ VI-4 Testing Procedure
On average, training took $\sim$90 min. After training, the order of conditions (DT, MM, and PM) was randomized. For each condition, subjects completed the tasks in the order of jar, then whiteboard, then plug. Subjects were given 3 and 1-minute remaining warnings. To minimize variance between the subjects, the placement of the target objects in the testing scene was kept consistent, and there were no distractor objects. Additionally, the jar and plug were modified to make the tasks slightly easier for novices: bright tape was added to the lid of the jar, and a socket adapter was used as the plug instead of an electrical cord so that the dynamics of the cable did not interfere with the task. Blue tape was also added to the adapter to make it easier to see. After attempting all of the tasks in a given condition, subjects filled out a questionnaire about their experience, measuring the system’s usability [3], the subject’s workload [9], and their feeling of presence in the remote environment. All questions were rated on a 7-point Likert scale.
@@ VII Results and Discussion
These results provide support for H1 and H2, indicating that the presented system can significantly improve novice operators’ ability to perform several tasks quickly and accurately. We also found that the predictive menu generally has a larger impact on both objective and subjective metrics than the manual menu, despite its relatively low accuracy of 60.on novice actions. We expect this impact to further increase as the number of possible actions and the accuracy of the predictor rise. The lack of support for H3 suggests that this menu system may preserve the operator’s sense of presence despite introducing additional visual elements not present in the remote environment; in fact, both MM and PM received higher average presence scores than DT. We attribute this to the minimally invasive nature of the hierarchical pie menu and affordances registered to the remote environment.
Our results show that contrary to conventional wisdom, designers of avatar robots need not choose between an immersive interface and using shared control: it is possible to achieve both in a single system. When integrating these two control paradigms, we suggest designers follow the philosophy presented here. For example, for shared control actions that reference the robot’s environment, directly overlaying visual elements corresponding to those actions onto the operator’s existing view lets the operator launch those actions while still focusing on their desired task. The manual menu presented here keeps the number of simultaneously presented icons low using a hierarchy, and this can be further improved for systems with large numbers of actions by using a predictive menu. With these techniques, designers of avatar systems can boost operator performance without sacrificing the benefits of an immersive direct control interface.

### v2

@@ VI Experiments
TABLE I: Differences between each interface across all tasks. , ∗∗, and ∗∗∗ denote $p\leq 0.05$, $p\leq 0.01$, and $p\leq 0.001$ respectively.
{tabu}
X[1.5] X X[r] X[r] X[r] X[r] X[r] Condition Success (%) ()($\uparrow$)( ↑ ) Time (s) ()($\downarrow$)( ↓ ) Usability ()($\uparrow$)( ↑ ) Workload ()($\downarrow$)( ↓ ) Presence ()($\uparrow$)( ↑ )
@@
Friedman p-value ∗∗∗0.0005 ∗∗0.0060 ∗∗0.0065 ∗∗∗0.0007 0.6004
Post-hoc p-value DT vs. MM ∗∗0.0066 0.0611 ∗∗0.0015 ∗∗0.0053 0.1308
DT vs. PM ∗∗∗0.0004 0.0115 ∗∗0.0061 ∗∗∗0.0004 0.5202
@@
Human subjects studies were conducted to evaluate differences between the DT, MM, and PM interfaces. All procedures were reviewed and approved by the UIUC IRB on Feb. 20, 2023. We formulated the following a priori hypotheses about the system:
@@
We recruited 20 student participants from the University of Illinois at Urbana-Champaign campus, 19 of whom completed the entire procedure. One participant requested to end the experiment during training due to nausea. Of the 19 participants, 11 were male, 7 were female, and one preferred not to say. Subjects were of age 19–32 (mean: 24) and self-reported their familiarity with robotics and controlling robots on average as 5.4 and 4.4 on a 7-point Likert scale [sarantakos2017social] respectively. None of the subjects had used TRINA before.
@@ VI-1 Basic Training
Subjects were trained to use the direct teleoperation interface and were introduced to several possible fault states. For example, if excessive force was applied to the arm, the subject would momentarily lose control of it. Subjects were given suggestions about how to resolve each of these faults. The assistive functionalities were demonstrated using the manual (MM) and predictive (PM) menus.
@@ VI-2 Task Introduction
Subjects were shown the three testing tasks and completed the tasks in-person to familiarize themselves with the specific features of the target objects. A researcher explained how task completion would be graded, and that subjects should try to complete tasks as quickly as possible with 5 min at most for each task. For the jar, the task was completed when the lid no longer was touching the jar body. For the whiteboard, the required writing was split into 19 segments and credit was given for each completed segment. For the plug, the task was completed when the subject had fully inserted the plug into the target socket.
@@ VI-3 Training Tasks
Subjects were coached through using the MM and PM on two training tasks which demonstrated each of the assistive actions in context. In the first task, a researcher handed TRINA a capped Expo marker, and the subject had to use TRINA to insert the tip of the marker into a square hole. Subjects were told to snap to the plane of the hole and turn off all rotational DoFs before inserting the marker into the hole. In the second task, subjects had to grasp and turn a dial for three full rotations. They were instructed to first snap to the circle of the dial, disable all but the x and roll DoFs to grasp the dial, and finally have only roll enabled to turn the dial.
@@ VI-4 Testing Procedure
On average, training took $\sim$90 min. After training, the order of conditions (DT, MM, and PM) was randomized. For each condition, subjects completed the tasks in the order of jar, whiteboard, then plug. Subjects were given 3 and 1 min remaining warnings. To minimize variance between the subjects, the placement of the target objects in the scene was kept consistent, and there were no distractor objects. Additionally, the jar and plug were modified to make the tasks slightly easier for novices: bright tape was added to the lid of the jar, and a socket adapter was used as the plug instead of an electrical cord. Blue tape was also added to the adapter to make it easier to see. After attempting all of the tasks in a given condition, subjects filled out a questionnaire about their experience, measuring the system’s usability [brooke1996sus], the subject’s workload [hart_development_1988], and their self-reported feeling of presence in the remote environment. All questions were rated on a 7-point Likert scale. Subjects would then immediately proceed to the next condition.
@@ VII Results and Discussion
These results provide support for H1 and H2, indicating that the presented system can significantly improve novice operators’ ability to perform several tasks quickly and accurately. We also found that the predictive menu generally has a larger impact on both objective and subjective metrics than the manual menu, despite its relatively low accuracy of 60.on novice actions. We expect this impact to further increase as the number of possible actions and the accuracy of the predictor rise. The lack of support for H3 suggests that this menu system preserves the operator’s sense of presence despite introducing non-physical visual elements; in fact, both MM and PM received higher average presence scores than DT. We attribute this to the minimally invasive nature of the hierarchical pie menu and affordances registered to the remote environment. We further found that both the MM and PM interfaces tend to increase the system’s usability and decrease the operator’s workload. Users can easily understand how to interact with both kinds of menus and use them to decrease the required cognitive effort to complete manipulation tasks.
Our results show that contrary to conventional wisdom, designers of avatar robots need not choose between an immersive interface and using shared control: it is possible to achieve both in a single system. When integrating these two control paradigms, we suggest designers follow the philosophy presented here. For example, for shared control actions that reference the robot’s environment, directly overlaying visual elements corresponding to those actions onto the operator’s existing view lets the operator launch those actions while still focusing on their desired task. The manual menu presented here keeps the number of simultaneously presented icons low using a hierarchy, and this can be further improved for systems with large numbers of actions by using a predictive menu.